Problem 8-29 Cullumber Toys management is considering eliminating product A, which has been showing a loss for several yearsRestate the income statement in segment margin format. TotalWhat would be the effect on income if product A were dropped? Net income would by $ LINK TO TEXT LINK TO VIDEO Management is
Answer #1
Segment Margin format Income Statement
A B C Total
Sales        2,244,000        1,407,000        1,813,400        5,464,400
Variable Expenses        1,684,000            601,000        1,099,500        3,384,500
Contribution Margin            560,000            806,000            713,900        2,079,900
Less: Traceable fixed expenses
Advertising            527,000            430,000            522,000        1,479,000
Depreciation              15,200              10,100              20,300              45,600
Segment Margin              17,800            365,900            171,600            555,300
Less: Common Expenses            277,900
Net Income            277,400
Net Income will reduce by 17800+15200 = $33,000
Units required to be sold = 560000/(10-4) = 93,333.33 units
i.e. 93,333 units